# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v2

@@ 3.2 Collective Predictive Coding and System 3
As a basis for the view of System 0/1/2/3, we introduce the CPC hypothesis. [136] hypothesized that the emergence of language can be understood through the lens of CPC. A visual representation of CPC is presented in Figure 4.
CPC expands the predictive coding concept [57] from individual cognition to collective societal systems. This framework allows the formalization of the emergence of symbol systems, including language, as a process of decentralized Bayesian inference [47, 143]. Essentially, language, viewed as a distribution of sentences, is derived from the posterior distribution of latent variables, given the distributed observations of multiple agents.
The basic form of CPC introduced by [136] can be mathematically expressed as
@@
where represents external (collective) representations is the set of representations by the -th agent, is the set of observations by the -th agent, is the set of sequence of actions taken by the -th agent, and is the sequence of rewards taken by the -th agent. The inference of corresponds to representation learning by the -th agent, and is assumed to be estimated through a language game within a group. Overall, symbol emergence by a group is performed to estimate in a decentralized manner. We assume that we cannot estimate the true posterior but can only estimate its approximate distribution .
Based on these premises, the variational free energy of CPC, which corresponds exactly to the model’s negative variational lower bound (ELBO), is defined as follows:
The inference of represents the representation learning process for the -th agent. The inference of is achieved through linguistic interactions within a group, such as naming games [143, 56, 64]. Collectively, these processes contribute to symbol emergence within a group, aiming to estimate in a decentralized manner.
@@
The subject of System 3 shifts the perspective from individual life to life as a collective entity. This can be understood by considering the question “Who is minimizing the free energy described in Equations 5-9?” The entity minimizing free energy is not the individual but rather the collective agents who share the variable defined in Equation 5, representing a collective form of life that engages in predictive coding of the world. While System 3’s operation corresponds to the dynamics of symbol and language emergence, one might assume that its learning relies solely on System 2 processes, such as individual language learning. However, we assert that System 3 is a collective emergent system. It adapts to its environment not only through the internal neural plasticity of individual agents but also by leveraging the plasticity of language itself in predictive coding. Consequently, the language formed through this process reflects collective predictive coding as a group.
@@ 3.3 Relationships and Dynamics Among Systems: A Cognitive and Temporal Perspective
| Consciousness | Non-conscious | Subconscious | Conscious | Supra-conscious |
@@
Temporal Scale and Model: First, the systems operate across distinct temporal scales, ranging from the super-fast physical and dynamical processes of System 0 to the super-slow emergent symbol systems of System 3. System 1, associated with fast processes, incorporates intuitive decision-making, whereas System 2, operating at slower speeds, is widely accepted in dual-process theory [69]. From the viewpoint of recent AI studies, Systems 1 and 2 correspond to world and language models [137, 33], respectively.
Representation and Language: The type of representation characterizes the nature of each system layer. System 0 automatically processes physical information without involving representations. System 1 handles implicit representations, such as non-linguistic features, which align with the concept of tacit knowledge [105]. System 2 employs explicit internal representations, corresponding to explicit knowledge. Although internal representations are not equivalent to language, they can be associated with inner speech, reflecting explicit internal processes. Finally, System 3 operates using external representations, corresponding to communication, such as outer speech between individuals.
@@ 4.1 Three Axes of Timescales
Bergson’s philosophy of time offers a broad perspective on diverse phenomena, ranging from matter to society, each operating on different timescales. When distinguishing between different these timescales, there are it is beneficial to differentiate between multiple axes that, which should not be conflated (see Figure 5).
@@ 4.2 Two Types of Memory and Thickness of the Present
Bergson’s theory of time explains natural phenomena through the combined effect of two types of “memory.” This distinction is often seen as a precursor to modern concepts of procedural and episodic memory [145, 106]; however, it has a broader scope. One type is “motor memory,” which refers to a series of motor patterns learned through repetition, such as skills and habits in humans. The other type is “pure memory,” which preserves original experiences before they become organized into habitual patterns. For example, through repeated trials of bicycle riding (r1, r2, r3), a motor memory for riding a bicycle (R) is acquired. However, one can also recall r1 as a specific life event on a particular date. Pure memory is what retains these unique instances.
@@
Let us now revisit the two types of recognition briefly mentioned in Section 2.4: a fast but limited automatic process and a slower but intentional process, the latter termed “attentive recognition” ([9], chap. 2). Both automatic and attentive recognition rely on motor memory and share the function of facilitating interactions within stereotyped environments. However, automatic recognition, operating with a “thin” window of the present, is confined to immediate motor responses. In contrast, attentive recognition, with its “thick” window, provides access to pure memory. This temporal opening allows active reconstruction of the recognized object as a mental image, and in some cases, even enables re-experiencing scenes from one’s personal past. Such temporal distinctions also likely apply to Systems 1 and 2.
@@ 4.3 Extended Definition of Motor Memory: Canalized Landscape
Bergson cites the example of the Sphex wasp stinging a caterpillar’s nine nerve centers in a precise order to paralyze but not kill it ([10], pp. 174–175). This intricate maneuver is neither learned nor reasoned but pre-configured. On an evolutionary scale, this has been achieved through countless variations and selections. Thus, behavioral optimization is completed even prior to an organism’s birth.
At both individual and evolutionary levels, establishing stable sensorimotor pathways in advance eliminates the need for a present window. In this sense, Bergson argues that as motor memory becomes perfected, it “steps outside of time”([9], p. 91). To illustrate this point further, we introduce the concept of the “Canalized Landscape.”
Bergson’s theory of perception involves the idea of knowledge outsourcing. In Matter and Memory, he rejects both representationalism and atomistic associationism, proposing a form of “direct realism” grounded in an organism’s potential actions. The idea that potential behavioral inducements are structured within the environment not only anticipates von Uexküll’s notion of the Umwelt [148] but also aligns remarkably well with the dynamical systems perspective discussed in Section 2.1. Each organism operates within its own action-inducing space, which extends beyond its body and is shaped through motor memory processes, including evolution and learning. Bergson likened its formation process to river formation over geological timescales, calling it “canalization” ([9], p. 280, [10], pp. 94-95). Following C. H. Waddington, a renowned geneticist who inherited this idea via A. N. Whitehead [107], we term the resultant space the “landscape” [149].
@@
If a river meanders suitably, water does not need to “think”; by flowing towards lower potential, it naturally follows the optimized “valley” paths. Similarly, the Sphex wasp’s behavior arises not solely from internal motor control but also from coupling with the caterpillar’s body shape and structure, allowing its actions to unfold automatically. Here, the spatial distinction between the internal and external becomes less significant. Instead, the focus is the temporally outsourcing of landscape formation (canalization) prior to the window of the present window.
@@ 4.4 Timescales of Canalization (System 0/1/2/3)
Finally, in humans who engage in social activities through symbols, canalization also occurs on the timescale of cultural evolution. In his fourth and final major work, The Two Sources of Morality and Religion, Bergson developed the concept of a socially canalized landscape. When people act collectively, certain behavioral channels are carved into the social landscape to resolve coordination problems, such as social norms, rules, and morality. Similar to evolutionary and individual learning landscapes, societal canalization fosters community-specific repertoires of behavior. Without this level of canalization, language usage would not be possible (System 3; see Section 2.3).
@@
First, Bergson reflects on the historical process by which the present window was gradually acquired through evolution. In simpler organisms, reactions “can then hardly be delayed” ([11], p. 22). Distant perception emerges when organisms become able to “defer the date of their fulfillment” ([11], p. 23) of responses. Thus, the presence of a present window and access to pure memory enables non-random improvisations in unforeseen situations, facilitating complex adaptive behaviors beyond what evolution or individual learning alone can achieve.
Moreover, on-the-fly variability promotes learning efficiency. Bergson notes: “thanks to this faculty, we have no need to await at the hands of chance the accidental repetition of the same situations, in order to organize into a habit concomitant movements; we make use of the fugitive image to construct a stable mechanism which takes its place” ([11], p. 98). Similarly, machine learning exhibits improved efficiency when supplemented by working memory [38, 146].
@@
“Our consciousness tells us that when we speak of our present we are thinking of a certain interval of duration. What duration? It is impossible to fix it exactly, as it is something rather elusive. My present, at this moment, is the sentence I am pronouncing. But [] my attention is something that can be made longer or shorter, like the interval between the two points of a compass.” ([12], pp. 178-179).
@@ 5 Discussion
5.1 Cultural Development and Creativity
Language is a fundamental aspect of culture. Research on the evolution of language and culture has often relied on the Iterated Learning Model (ILM). This model simulates the intergenerational transmission of language, modeling how parental language production and child language acquisition form a cycle. This explains how compositionality, a universal feature of human language, emerges through cultural evolution. This phenomenon has been demonstrated through agent-based simulations [74, 75, 78], mathematical models [15, 40, 77], and laboratory-based language evolution experiments (experimental semiotics) [76, 78, 115]. The ILM was designed to capture the propagation of expressions. However, it fails to adequately model the interaction between expressions and environmental adaptation, which involves physical interactions and (internal) representation learning, such as the generation of meaning for these expressions. By contrast, EmCom models focus on the process of forming external representations within a developmental framework. Building on this, Generative EmCom integrates generative model-based representation learning to establish a novel theoretical framework [141]. This framework, formalized as CPC, extends beyond language and encompasses a wide range of cultural expressions.
@@ 5.1 Cultural Development and Creativity
Artistic activities, such as music, literature, and performance, provide valuable avenues for exploration within CPC and the System 0/1/2/3 framework. Recent advancements in generative AI technologies, such as diffusion models for image creation and music generation techniques, have sparked debates regarding whether AI can truly possess creativity [55, 111, 59]. While much of this discourse focuses on the “generation” aspect, the essence of creativity extends beyond production to include “evaluation” and “acceptance,” as creative works are deeply rooted in human emotions and embodied sensory experiences. From a systemic perspective, creativity emerges through a dynamic interplay of various components, as emphasized by Csikszentmihalyi”s system model of creativity [21]. Creativity is not solely inherent in the artifact but arises from interactions between the creator, artifact, evaluating community, and contextual factors shaping its creation, such as cultural trends, historical events, and sociopolitical influences. These elements form a cyclical system in which each part continually affects and is affected by others; the artist draws inspiration from the community and context, while the community evolves its standards through the artist”s contributions.
@@
Modeling such intricate dynamics within CPC and the System 0/1/2/3 framework poses a significant challenge for AI because it requires the integration of generative, evaluative, and emotional components to capture the reciprocal and emergent nature of human creativity. Saunders et al. compiled a review of computational models of creativity that emphasize social interactions in the context of ALife. They also argued for the importance of studying creativity as a system theory [113]. From a perspective of emotional feeling construction, discussions have emerged regarding the role of predictive coding in the sensation of interoceptions [116, 3]. By replacing exteroception, e.g., visuals and haptics, of sensory observations in CPC with sensations of interoceptions, artistic or entertaining expressions, including music, poetry, and performance, will fall within the scope of CPC, and therefore, within the scope of the System 0/1/2/3 framework. Considering the parallelism between music and language, the systematic correspondence between the emergence of two different symbol systems, i.e, music and language, has been discussed [135].
@@ 5.2 Consciousness and its collectiveness
Consciousness and unconsciousness also have a collective aspect, as exemplified by Jung’s concept of collective unconsciousness [68]. The collective dimension of consciousness is a critical factor in understanding the human mind. Furthermore, Lacan highlighted the influence of language on the unconscious within the context of structuralism, emphasizing the role of linguistic structures in shaping the unconscious mind [79].
Recent research has also explored the relationship between CPC and the structure of qualia [139]. Building on System 0, the learning of representations through Systems 1 and 2 and the linguistic interactions facilitated by System 3 raise profound questions about how consciousness interacts with the structure of language formed in society, i.e., System 3. These interactions remain an open area of inquiry, particularly regarding how language and representation processes affect conscious and unconscious states.

### v3

@@ 3.2 Collective Predictive Coding and System 3
As a basis for the view of System 0/1/2/3, we introduce the CPC hypothesis. taniguchi2024collective hypothesized that the emergence of language can be understood through the lens of CPC. A visual representation of CPC is presented in Figure 4.
CPC expands the predictive coding concept (hohwy2013predictive) from individual cognition to collective societal systems. This framework allows the formalization of the emergence of symbol systems, including language, as a process of decentralized Bayesian inference (hagiwara2019symbol; taniguchi2022emergent). Essentially, language, viewed as a distribution of sentences, is derived from the posterior distribution of latent variables, given the distributed observations of multiple agents.
The basic form of CPC introduced by taniguchi2024collective can be mathematically expressed as
@@
[…]
The inference of represents the representation learning process for the -th agent. The inference of is achieved through linguistic interactions within a group, such as naming games (taniguchi2022emergent; hoang2024emergent; inukai2023recursive). Collectively, these processes contribute to symbol emergence within a group, aiming to estimate in a decentralized manner.
It is crucial to clarify that the FEP is assumed to operate hierarchically across all systems, starting from the sensorimotor level. The observations, o, and actions, a, in the CPC formulation represent precisely these sensorimotor signals. These signals are generated through the physical interactions of System 0 with the environment and are subsequently processed by System 1 and above, forming a continuous perception-action loop governed by the imperative to minimize free energy.
@@
[…]
While System 3’s operation corresponds to the dynamics of symbol and language emergence, one might assume that its learning relies solely on System 2 processes, such as individual language learning. However, we assert that System 3 is a collective emergent system. It adapts to its environment not only through the internal neural plasticity of individual agents but also by leveraging the plasticity of language itself in predictive coding. Consequently, the language formed through this process reflects collective predictive coding as a group.
@@ 3.3 Relationships and Dynamics Among Systems: A Cognitive and Temporal Perspective
[…]
@@
Temporal Scale and Model: First, the systems operate across distinct temporal scales, ranging from the super-fast physical and dynamical processes of System 0 to the super-slow emergent symbol systems of System 3. System 1, associated with fast processes, incorporates intuitive decision-making, whereas System 2, operating at slower speeds, is widely accepted in dual-process theory (kahneman2011thinking). From the viewpoint of recent AI studies, Systems 1 and 2 correspond to world and language models (taniguchi2023world; friston2021world), respectively.
Representation and Language: The type of representation characterizes the nature of each system layer. System 0 automatically processes physical information without involving representations. System 1 handles implicit representations, such as non-linguistic features, which align with the concept of tacit knowledge (polanyi1966tacit). System 2 employs explicit internal representations, corresponding to explicit knowledge. Although internal representations are not equivalent to language, they can be associated with inner speech, reflecting explicit internal processes. Finally, System 3 operates using external representations, corresponding to communication, such as outer speech between individuals.
@@
Cognitive Processes and Conscious Processing: Each layer is characterized by distinct cognitive processes, as summarized in Table 1. System 0 operates reflexively, without any form of consciousness. The distinction in the table regarding System 1 (“Subconscious”) and System 2 (“Conscious”) refers primarily to the concept of conscious information processing, common in dual-process theories. In this view, System 1 manages tasks automatically, without conscious deliberation, whereas System 2 involves explicit, conscious processing.
This discussion of information processing, however, must be distinguished from the broader issue of consciousness as subjective experience, which includes phenomena like qualia. In that regard, we do not claim that consciousness is exclusive to System 2. Rather, we propose that the fast, sensory-level processing of System 1 provides the essential raw material for phenomenal consciousness. This hierarchical view is well-aligned with Bergson’s philosophy of time. He distinguished between foundational “sensory qualities” (qualités sensibles), which is the equivalent of modern qualia, arising at the minimal temporal resolution of perception, and their integration into a thicker “duration” (durée) where meaningful, flowing experience is formed. Our framework maps onto this hierarchy: System 1 is associated with the emergence of these foundational sensory qualities, while System 2’s reflective consciousness operates within this wider "duration," acting upon the materials provided by System 1.
System 3, in turn, invites speculation about collective cognitive phenomena. The notion of cognitive phenomena that transcend the individual has a long history in various schools of thought, exemplified by concepts like Durkheim’s “collective consciousness” or Jung’s “collective unconscious” (durkheim1893division; jung1969archetypes). Our framework, with its specific focus on temporality, allows us to pose a related but distinct question: could the vast temporal difference between individual cognition and the super-slow evolution of societal symbol systems give rise to a form of consciousness? While this remains a profound open question beyond the scope of this paper, we use the term "supra-conscious" to denote this hypothetical possibility—a collective cognitive phenomenon that is yet to be explored.
@@ 4.1 Three Axes of Timescales
[…]
@@ 4.2 Two Types of Memory and Thickness of the Present
[…]
@@
Let us now revisit the two types of recognition briefly mentioned in Section 2.4: a fast but limited automatic process and a slower but intentional process, the latter termed “attentive recognition” (bergson1896, chap. 2). Both automatic and attentive recognition rely on motor memory and share the function of facilitating interactions within stereotyped environments. However, automatic recognition, operating with a “thin” window of the present, is confined to immediate motor responses. In contrast, attentive recognition, with its “thick” window, provides access to pure memory. This temporal opening allows active reconstruction of the recognized object as a mental image, and in some cases, even enables re-experiencing scenes from one’s personal past. Such temporal distinctions also likely apply to Systems 1 and 2.
@@
Husserl, James, and Bergson—writing almost concurrently—converged on the idea that a bare, given succession or long-term memory cannot account for the flow experience of time; a distinct structure of time-consciousness is required. James and Bergson locate this structure in an extended but finite scope of the present, wide enough to encompass the lived flow of time, whereas Husserl’s retention–protention schema leaves that width formally unbounded; this classical contrast re-emerges today as the debate between retentional and extensional models of temporal experience (Dainton2023). Bergson goes further, treating the span of the present as an evolutionary variable: as it lengthens, organisms can integrate a broader sweep of ongoing sensory change within a single present, thereby enabling ever-richer forms of cognition (Winkler2006).
@@ 4.3 Extended Definition of Motor Memory: Canalized Landscape
Bergson cites the example of the Sphex wasp stinging a caterpillar’s nine nerve centers in a precise order to paralyze but not kill it (bergson1907, pp. 174–175). This intricate maneuver is neither learned nor reasoned but pre-configured. On an evolutionary scale, this has been achieved through countless variations and selections. Thus, behavioral optimization is completed even prior to an organism’s birth.
At both individual and evolutionary levels, establishing stable sensorimotor pathways in advance eliminates the need for a present window. In this sense, Bergson argues that as motor memory becomes perfected, it “steps outside of time”(bergson1896, p. 91). To illustrate this point further, we introduce the concept of the “Canalized Landscape.”
[…]
@@
[…]
By slicing cognition along temporal rather than spatial inside/outside lines, we obtain a staged and inclusive picture that neither depends wholly on representation nor banishes it. Treating representations as timescale–contingent strategies—emerging only when slower loops become relevant—avoids the rigidity of any single “-ism” and embodies the flexible stance (Constant2021). This temporal–slicing scenario allows for a dynamic blending of enaction and representation in proportion to the available timescale: shorter loops favour direct sensorimotor engagement, whereas longer loops progressively introduce representational strategies. Bergson had already anticipated this logic: by focusing on how evolutionary elongation of an organism’s temporal architecture—the delays built into sensorimotor coupling—creates room for “memory–images,” he explained the rise of attentive recognition and interpretive acts. In this sense, his view converges with the contemporary account of representations as timescale-dependent emergents (Bruineberg2014).
@@ 4.4 Timescales of Canalization (System 0/1/2/3)
Finally, in humans who engage in social activities through symbols, canalization also occurs on the timescale of cultural evolution. In his fourth and final major work, The Two Sources of Morality and Religion, Bergson developed the concept of a socially canalized landscape (Hirai2025). When people act collectively, certain behavioral channels are carved into the social landscape to resolve coordination problems, such as social norms, rules, and morality. Similar to evolutionary and individual learning landscapes, societal canalization fosters community-specific repertoires of behavior. Without this level of canalization, language usage would not be possible (System 3; see Section 2.3).
@@
First, Bergson reflects on the historical process by which the present window was gradually acquired through evolution. In simpler organisms, reactions “can then hardly be delayed” (bergson1911, p. 22). Distant perception emerges when organisms become able to “defer the date of their fulfillment” (bergson1911, p. 23) of responses. Thus, the presence of a present window and access to pure memory enables non-random improvisations in unforeseen situations, facilitating complex adaptive behaviors beyond what evolution or individual learning alone can achieve.
Moreover, on-the-fly variability promotes learning efficiency. Bergson notes: “thanks to this faculty, we have no need to await at the hands of chance the accidental repetition of the same situations, in order to organize into a habit concomitant movements; we make use of the fugitive image to construct a stable mechanism which takes its place” (bergson1911, p. 98). Similarly, machine learning exhibits improved efficiency when supplemented by working memory (graves2016; vinyals2016).
@@
“Our consciousness tells us that when we speak of our present we are thinking of a certain interval of duration. What duration? It is impossible to fix it exactly, as it is something rather elusive. My present, at this moment, is the sentence I am pronouncing. But [] my attention is something that can be made longer or shorter, like the interval between the two points of a compass.” (bergson1946, pp. 178-179).
@@ 5 Discussion
[…]
@@
Artistic activities, such as music, literature, and performance, provide valuable avenues for exploration within CPC and the System 0/1/2/3 framework. Recent advancements in generative AI technologies, such as diffusion models for image creation and music generation techniques, have sparked debates regarding whether AI can truly possess creativity (ho2020denoising; rombach2022high; huang2024symbolic). While much of this discourse focuses on the “generation” aspect, the essence of creativity extends beyond production to include “evaluation” and “acceptance,” as creative works are deeply rooted in human emotions and embodied sensory experiences. From a systemic perspective, creativity emerges through a dynamic interplay of various components, as emphasized by Csikszentmihalyi”s system model of creativity (csikszentmihalyi2015systems). Creativity is not solely inherent in the artifact but arises from interactions between the creator, artifact, evaluating community, and contextual factors shaping its creation, such as cultural trends, historical events, and sociopolitical influences. These elements form a cyclical system in which each part continually affects and is affected by others; the artist draws inspiration from the community and context, while the community evolves its standards through the artist”s contributions.
@@
Modeling such intricate dynamics within CPC and the System 0/1/2/3 framework poses a significant challenge for AI because it requires the integration of generative, evaluative, and emotional components to capture the reciprocal and emergent nature of human creativity. Saunders et al. compiled a review of computational models of creativity that emphasize social interactions in the context of ALife. They also argued for the importance of studying creativity as a system theory (saunders2015computational). From a perspective of emotional feeling construction, discussions have emerged regarding the role of predictive coding in the sensation of interoceptions (seth2013interoceptive; barrett2015interoceptive). By replacing exteroception, e.g., visuals and haptics, of sensory observations in CPC with sensations of interoceptions, artistic or entertaining expressions, including music, poetry, and performance, will fall within the scope of CPC, and therefore, within the scope of the System 0/1/2/3 framework. Considering the parallelism between music and language, the systematic correspondence between the emergence of two different symbol systems, i.e, music and language, has been discussed (taniguchi2021parallelism).
Our full System 0/1/2/3 framework also leads to novel, falsifiable predictions for future research. First, it predicts that a group’s shared symbol system will be systematically shaped by their physical embodiment (System 0). This could be tested in a robotics experiment comparing two groups of robots with different morphologies (e.g., a rigid gripper vs. a soft manipulator) to see if their emergent languages diverge to reflect their differing sensorimotor affordances. Second, the framework predicts that collective adaptation to a new environment will be more efficient through “semiotic plasticity” (re-negotiating the shared language in System 3) than through only individual learning (in Systems 1/2). This can be tested in agent-based simulations by comparing a group that can modify its language to one with a frozen language, measuring the speed of adaptation after an environmental shift.
@@ 5.2 Consciousness and its collectiveness
Consciousness and unconsciousness also have a collective aspect, as exemplified by Jung’s concept of collective unconsciousness (jung1969archetypes). The collective dimension of consciousness is a critical factor in understanding the human mind. Furthermore, Lacan highlighted the influence of language on the unconscious within the context of structuralism, emphasizing the role of linguistic structures in shaping the unconscious mind (lacan2006ecrits).
Recent research has also explored the relationship between CPC and the structure of qualia (taniguchi2024constructive). Building on System 0, the learning of representations through Systems 1 and 2 and the linguistic interactions facilitated by System 3 raise profound questions about how consciousness interacts with the structure of language formed in society, i.e., System 3. These interactions remain an open area of inquiry, particularly regarding how language and representation processes affect conscious and unconscious states.